OpenStudy (anonymous):

x^2=-2x+33 factor show steps

OpenStudy (anonymous):

would it be easier to solve graphically?

OpenStudy (solomonzelman):

I meant completing the square.

OpenStudy (anonymous):

but factoring is what im needing to do

OpenStudy (solomonzelman):

wait, I don't you can factor it

OpenStudy (solomonzelman):

not into integers at least

OpenStudy (solomonzelman):

completing the square I find ideal.

OpenStudy (solomonzelman):

1) add 2x to both sides 2) add 1 to both sides

OpenStudy (anonymous):

ok i have to go acually srry thanks for ur imput

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x^2=-2x+33 }\) \(\large\color{black}{ \displaystyle x^2\color{blue}{+2x}=-2x+33\color{blue}{+2x} }\) \(\large\color{black}{ \displaystyle x^2+2x=33 }\) \(\large\color{black}{ \displaystyle x^2+2x\color{blue}{+1}=33 \color{blue}{+1} }\) \(\large\color{black}{ \displaystyle x^2+2x+1=34 }\)

OpenStudy (solomonzelman):

now, complete the square (the left side is a perfect square trinomial) if you haven't yet learned to complete the square, we can use the quadratic formula